OpenStudy (anonymous):

What is the first step in solving 4(x + 2) = 10 + x? A. Distribute by multiplying 4 to the terms in the parentheses. B. Multiply by 4 on both sides. C. Subtract 2 on both sides. D. Subtract x on both sides.

Parth (parthkohli):

What do you think will it be?

Parth (parthkohli):

It's the multiplication of 4 and (x + 2). PEMDAS says that multiplication should be done first. So, you should apply distributivity.

Parth (parthkohli):

Distributivity says: \(\Large \color{MidnightBlue}{\rightarrow a(x + y) = ax + ay }\)

Parth (parthkohli):

So, 4(x + 2) = 4x + 4(2)

OpenStudy (anonymous):

4x+8=10+x If you're solving for x, you would subtract 8 to both sides and subtract x from both sides 3x=2 Divide by 3 x=2/3 Parenthesis > Exponent > Multiplication/Division > Addition/Subtraction Answer is A :D

OpenStudy (anonymous):

Thanks everyone, I was going to say A but you already said it first :(

OpenStudy (anonymous):

Welcome